Aniyah2:

Ally surveyed 5 classrooms. She wanted to know the number of students in each class who love painting. The data she collected from the 5 classrooms is below. 10, 7, 9, 11, 8

KayvonGC:

9

Aniyah2:

@kayvongc wrote:
9
ty-,-

steevh:

9

Aniyah2:

@steevh wrote:
9
lol ty

steevh:

total 45/5(classrooms)

KayvonGC:

im just smart like that

Aniyah2:

@kayvongc wrote:
im just smart like that
sike

KayvonGC:

not funny

Aniyah2:

@kayvongc wrote:
not funny
very

KayvonGC:

........

steevh:

lol

KayvonGC:

hi sam

sam1235:

lol hi kayvon

Fortish:

Hi, please don't give the user the direct answer. Guide the user to it.

Fortish:

@fortish wrote:
Hi, please don't give the user the direct answer. Guide the user to it.
Its against the rules, please and ty.

Fortish:

@kayvongc wrote:
im just smart like that
@sam1235 wrote:
lol hi kayvon
@kayvongc wrote:
hi sam
@steevh wrote:
lol
Responses like these are considered spam also against the rules, please read the rules if you didn't.

KayvonGC:

good for u ig nerd

Fortish:

@kayvongc wrote:
good for u ig nerd
I won't report ya'll but just keep that in mind. @aniyah2 please close question

KayvonGC:

please go touch grass

KayvonGC:

thx bye

sam1235:

whats going on?

KayvonGC:

idk

Fortish:

@shadow please lock this question

KayvonGC:

or what

Aniyah2:

KayvonGC:

not funny

Aniyah2:

Who said it was-

KayvonGC:

who are u talking to

TutorJohnfoster:

The value that could be the best measure of the center of the data is the mean. mean=\[\frac{Total\space of all\space items }{ Number\space of items }\] \[\frac{ 10+7+9+11+8 }{ 5 }\] \[\frac{ 45 }{ 5 }\] 9